OpenStudy (anonymous):

What is the equation in point−slope form of the line passing through (−3, −4) and (0, 2)? (y − 8) = 2(x − 3) (y + 3) = −2(x + 4) (y − 2) = −2(x + 4) (y + 4) = 2(x + 3)

OpenStudy (anonymous):

@zepdrix long time no talk zep :( how ya been? you busy rn?

OpenStudy (amity):

M=y2-y1/x2-X1,y-y1=m(x-x1)

OpenStudy (amity):

M=y2-y1/x2-X1,y-y1=m(x-x1)

RhondaSommer (rhondasommer):

do you know how to get slope?

OpenStudy (anonymous):

i know the forms but when i did it, i ended up with a fraction for m

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you find the slope first but i ended up with a fraction. so i then did point-slope and i ended up with a weird number that wasnt even part of the answers

TheSmartOne (thesmartone):

\(\sf\Large Slope = \frac{\color{blue}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{purple}{x_1}}\) Where the two points are in the form: \(\sf\Large (\color{purple}{x_1},\color{red}{y_1})~and~(\color{green}{x_2},\color{blue}{y_2})\)

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

i did that lol

OpenStudy (anonymous):

my problem is that when i did it, i got 5/4

OpenStudy (anonymous):

for m

OpenStudy (amity):

M=2+4/0+3=2

TheSmartOne (thesmartone):

show me your work, you make a mistake

OpenStudy (anonymous):

lemme do it on paper

OpenStudy (amity):

M=2+4/0+3=2

OpenStudy (anonymous):

ok so then you put it into point slop form right?

OpenStudy (amity):

Last option is correct

OpenStudy (amity):

Last option is correct

RhondaSommer (rhondasommer):

yes you place which ever set of points in you want

OpenStudy (anonymous):

thanks a mil guys

TheSmartOne (thesmartone):

Anytime! :D

OpenStudy (amity):

Welcome

OpenStudy (amity):

Welcome